OpenStudy (anonymous):

determine the volume of a 0.530 M KMno4 solution required to completely reast with 2.6 g of zn

OpenStudy (anonymous):

i try and i got 75.02 mL . but wrong. is it I doing wrong?

OpenStudy (anonymous):

Can u write the equations?

OpenStudy (anonymous):

I mean just the ionic equations...

OpenStudy (anonymous):

Ok see since nothing is mentioned i assume the reaction is carried out in a neutral medium, that means Mn reduces it state from +7 to +4. And Zinc is oxidized from 0 to +2. ok?

OpenStudy (anonymous):

Given is 2Mno4- + 8H+ + 5Zn -> 2Mn2+ + H2O +5Zn2+ But i dont see K there, so i dont know if i can use this..

OpenStudy (anonymous):

Ah ok so now just observe the coeffecients of Mn and Zn in the reactants the rest is useless. Mn coeffecient-2. Zn-5. Do you know law of equivalents? M1V1e1=M2V2e2

OpenStudy (anonymous):

yes.. i knew it.. im using that just know, but got the wrong answer.

OpenStudy (anonymous):

Well actually in this reaction which is a redox reactions only the ions which undergo a change in oxidation state matter.If you see Mn on the left you have it in +7 oxidation state but on the right it is in +2 so e1=5. Now Zn changes from 0 to +2 so its e2=2. Now M2 V2 of zinc is nothing but the number of moles of Zn=2.6/65.38=0.04 So 5*x*0.53=0.04*2. You get x as 0.0301L which is 30.1mL

OpenStudy (anonymous):

M1V1=n2 .. ok, i got it..